OpenStudy (anonymous):

find the point on the curve r(t)=(sin(3t))i+(cos(3t))j+(4t)k at a distance 5pi/12 units along the curve from the point (0,-1,4pi/3) in the direction of increasing arc length

OpenStudy (anonymous):

i need verification..

OpenStudy (anonymous):

r'(t)=(3cos(3t))i-(3sin(3t))j+4k |r'(t)|=5 S=\[\int\limits_{t(0)}^{t}5d \tau\]

OpenStudy (anonymous):

and t(0)= pi/3 so that it satifies original point (0,-1,4pi/3)

OpenStudy (anonymous):

i think \[5\tau-5(\pi/3)\] has to be set to equal 5pi/12..?

OpenStudy (mathmate):

The integral is correct, but it should be integrated with respect to t. After integration, it would be 5t] from pi/3 to t1 = (5/12)pi Solve for t1.

OpenStudy (anonymous):

right, i just didn't want to confuse myself so just added another variable. so after solving, i get 5pi/12 which then gets plugged into r(5pi/12)=(sin(3*5pi/12))i+(cos(3*5pi/12))j+4(5pi/12)k...?

OpenStudy (mathmate):

All good, great job!

OpenStudy (anonymous):

ok great, finally got something right, thank you

OpenStudy (mathmate):

You're welcome!